a coil with 50 turns and area 10cm2 is oriented with a plane perpendicular with a 0.75-T magnetic field. if the coil is rotated over(rotated through 180) in 0.20s. what is the average emf induced in it. 0.375v

1
Expert's answer
2022-06-15T07:03:35-0400

Given:

"N=50; \space A=10cm^{2}=10^{-3}m^{2}; \space B=0.75T; \space t=0.2s;"

"\alpha=180^{0};"

Solution:

Formula of induced EMF:

"E=-N\frac{\Delta\phi}{\Delta t};" Let's find "\Delta\phi" - magnetic flux though the coil;

"\Delta\phi==\int_0^αB\delta(Acosα)=\int_0^\pi B\delta(Acosα)=-2BA;" here "B\space and \space A \space constant;"

"E=-N\frac{\Delta\phi}{\Delta t}=N\frac{2BA}{t}=50\times\frac{2\times10^{-3}m^{2}\times0.75T}{0.2s}=0.375V;"

Answer:

"E=0.375V."